OpenStudy (anonymous):

21. Write the equation of a line that is perpendicular to the given line and that passes through the given point. y - 3 = 8/3(x + 2); (-2, 3)

OpenStudy (anonymous):

i think it might be (0,-2)

OpenStudy (anonymous):

What is the slope in the first formula?

OpenStudy (anonymous):

Basic equation of the line : y=mx+b, where m is a slope If two lines are perpendicular it means that if one has slope m1 then the other one has m2=-1/m1 (negative reciprocal) Equation of the line passing through point (x1,y1): y – y1 = m(x – x1) 1. find the slope of line 2 if line1 y=3x+2 has slope m1=3, then slope of line2 m2=-1/m1=-1/3 2.substitute m2,(x1, y1)=(1,-1) y-y1=m(x-x1) y-(-1)=-1/3(x-1) y+1=-1/3x+1/3 3.write equation in basic form y=-1/3x-2/3---answer :D HOPE THIS HELPS!!